OpenStudy (scorcher219396):

There are two tangent lines to the parabola y=x^2 that pass through the point (0,-4). Find the coordinates of the points where these tangent lines intersect the parabola.

OpenStudy (scorcher219396):

I know f'(x) is 2x... and i know point slope formula... but I have no idea what to do... @Kainui

OpenStudy (scorcher219396):

Except just trying points and finding the formula for a tangent line which has a y intercept of -4, but that seems inefficient and not what I'm supposed to be doing

OpenStudy (science0229):

How about this. Set the equation of the tangent line as y=mx-4

OpenStudy (scorcher219396):

Ok

OpenStudy (science0229):

Now, solve the system of equations. y=mx+b y=x^2

OpenStudy (science0229):

You would get\[x^2=mx-4 \rightarrow x^2-mx-4=0\]

OpenStudy (scorcher219396):

Wouldn't it be x^2-mx+4?

OpenStudy (science0229):

Oops Yeah. My bad.

OpenStudy (science0229):

\[x^2-mx+4=0\]

OpenStudy (science0229):

Now, because y=mx-4 and y=x^2 are tangent of each other, how many intersections do they have?

OpenStudy (scorcher219396):

A single point But two, one on each side of the parabola

OpenStudy (science0229):

For a quadratic equation, if it has one solution, its discriminant has to be 0, right?

OpenStudy (scorcher219396):

Right

OpenStudy (science0229):

What is the discriminant for x^2-mx+4=0

OpenStudy (scorcher219396):

Actually, can't you plug in 2x there for the m in x^2-mx+4, which will give you the x coordinates of the tangent points? Then plug those in to get the y's?

OpenStudy (science0229):

You can't do that. The reason is 2x represents the slope of a tangent line of an ARBITRARY point on y=x^2 Because there are only 2 points where these tangent lines intersect the parabola, you can't do that.

OpenStudy (scorcher219396):

Right, but that arbitrary point is the same arbitrary point that's being solved for with the x in the substituted formula from the system which is solving for that point

OpenStudy (scorcher219396):

I tried it and got the right answer, +-2,4

OpenStudy (science0229):

m should be 4, -4

OpenStudy (science0229):

I worded it wrong. It's not arbitrary. 2x means for ANY point on y=x^2

OpenStudy (science0229):

x^2-mx-4=0 doesn't hold true for ANY points on y=x^2

OpenStudy (scorcher219396):

But that m is there because it was originally part of the equation for the slope of the tangent line, which by definition, has the same slope as the tangent point on the parabola

OpenStudy (science0229):

m was meant to be the slope of the tangent lines THAT PASSES THROUGH (0,4)

OpenStudy (scorcher219396):

The line that passes through 0,4 that we are forcing to have the same slope as the parabola at some point x in order to be tangent, therefore has a slope of 2x as well

OpenStudy (science0229):

I agree. But, you have to understand that x is a variable, meaning that it changes.

OpenStudy (science0229):

The tangent point doesn't change, meaning that it has to stay at one place.

OpenStudy (scorcher219396):

But it still represents the same point We just don't know what that point is

OpenStudy (scorcher219396):

And therefore are using x to represent it

OpenStudy (science0229):

We already used 'x' in y=x^2 You can't use 'x' again.

OpenStudy (science0229):

'x' is the variable that is used in functions.

OpenStudy (science0229):

The correct way of doing this your way would be this:

OpenStudy (science0229):

I intentionally represent only one tangent line for my own conveniance.